a small square is constructed. Then a new square is made by increasing each side by 2 meters. The perimeter of the new square is 3 meters shorter than five times the length of one side of the original square. find the dimensions of the original square.
Answer by ewatrrr(24785)


Hi
Let x represent the length of a side in small square. P = 4x
P(larger square) = 4(x+2m) = 4x + 8m
Question states***
4x + 8m -3m = 5x
Solving for x
5m = x, the length of a side in small square (the original one)
CHECKING our Answer***
4*7 - 3 = 28 -3 = 25 =5*5